GreenHab Report – December 3rd

GreenHab Report

Trisha Randazzo

December 3, 2017

Sol 2

Environmental control: (Choose which is appropriate and explain further if needed)

Ambient (no heating or/cooling)

Ambient with window/door open

·         Windy so unable to keep door open

Heating

·         Functioning nominally

Cooling

·         Not availabloe

Both heating and cooling

·         Nominal

Shade cloth on/off

·         On

Average temperature: (N/A until new sensor is delivered)

·         N/A

Changes to crops:  Note all emergence of seedlings, death of seedlings, etc

·         One tomato is doing poorly, but everyone else is growing quite well!!

Daily water usage for crops:

·         Complete with no issues

Time(s) of watering for crops:

·         12:55

Morning research observations: (Currently not operational)

·         N/A

Changes to research plants:

·         N/A

Daily watering and amount of water used:

·         2 gallons of all soils

Aquaponics:  (Currently not operational)

·         N/A

Narrative:  Any other information you want to share

·         Planted crew gifts of paper-whites. Two containers in the Green-Hab, one container in the Hab.

Support/supplies needed:

·         Awaiting:

Operations Report – December 3rd

Crew 184 Operations Report  03/12/2017

SOL: SOL 2

Name of person filing report: HuntJ

Non-nominal systems: Toilet Tank, Water (static)

Notes on non-nominal systems:  Toilet Tank: Tank was particularly filled up during purge and caused all crew to have to move upstairs for duration of purge. Workaround is to purge every other day of ops (e.g Sol 4 is next purge) to make upkeep easier on all.

We also request a resupply of the TB-Cide Quat Cleaner Deordorizer/Disinfectant. We are out after the last purge.

 

                                                    Water (Static): Water is at 50 Gallons which is alarmingly low. Our trailer water is at 1000 gallons but we need the means to transfer the water. Request to get Trailer water moved to Static water or permission and high level instructions to do it ourselves.

Generator (hours run): 13.2 Hours

Solar— SOC % (Before generator is run at night) Not nightfall as of submission, but will report this info in the next report.

Diesel –  50% full

Propane –  Did not check (No EVA), assume to be 79 percent volume due to past readings and trends.

Ethanol Free Gasoline (5 Gallon containers for ATV) –   gallons. 5 gallons

Water (trailer) –   1000 Gallons

Water (static) –   50 Gallons

Trailer to Static Pump used –  No

Water (loft) – Static to Loft Pump used –  Yes

Water Meter:  20 gallons (46 once Pump used)

Toilet tank emptied: Yes

ATV’s Used: None

Oil Added? No

ATV Fuel Used:  None

# Hours the ATVs were Used today: 00

Notes on ATVs: N/A

Deimos rover used: Not used

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used:  ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used:  Not used

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used:  Not Used

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used:  Not Used

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where?  No

General notes and comments:

Summary of internet:  Still good so far.

Summary of suits and radios: All nominal

Summary of Hab operations: Static Water Tank almost empty, need means to transfer Trailer water over to Static.

Summary of GreenHab operations:  All nominal

Summary of ScienceDome operations: All nominal

Summary of RAMM operations: Not Operational

Summary of health and safety issues: All nominal

Questions, concerns and requests to Mission Support: Is there a way we can check to see how much internet we are using during the day? We tried the link provided in a different log template but it didn’t take us to anything.

Sol Summary – December 3rd

Crew 184 Daily Summary Report  03 Dec 2017

MDRS Daily Summary Report for sol 2

Summary Title:  Getting settled at MDRS!

Mission Status:  Today has been a busy day preparing both ourselves and the MDRS for the rest of our mission.  We spoiled ourselves with a hearty breakfast today (and to get rid of some of our fresh food!).

Today has been one of excitement preparing for our first EVA tomorrow.  We have encountered our first mission failure!

During the high winds today we heard a bang as the outer airlock door was blown open.  The inner airlock door held closed and thus we didn’t all die!  We immediately closed the door to the EVA Prep room to form a second seal to the martian atmosphere.  During tomorrows familiarization EVA we will enter the EVA Prep room, close the EVA Prep room door, and use that as a depressurization chamber.  Immediately upon departure on tomorrows EVA we will repair the outer airlock door.

We have discovered today that all the planning in the world still requires real-time flexibility as numerous activities took much longer than we expected.  For example the Exercise Bike took several hours and modifications to assemble and get working correctly, much more than we had allocated.  That being said, we have successfully completed all objectives and are ready to begin completing our science objectives tomorrow!.

Sol Activity Summary:

1.    Inventoried and organized the MDRS Hab and Greenhab

2.    Completed all reports

a.    Greenhab

b.    HSO

c.    Mission Plan

d.    Operations Report

e.    Sol Summary

f.     EVA Request

g.    Journalist

3.    Assembled exercise equipment (Stationary Bike / Weights)

4.    Got more detailed training from previous crewmembers

5.    Crew Exercise

6.    Internet troubleshooting (WiFi modem nonfunctional – connection working straight to modem)

7.    Cleaning and organizing MDRS

8.    Radio check with all radios (all functional except for 1)

9.    Communication check at all MDRS buildings with Habitat

10. Neuroblock Feasability Training

Look Ahead Plan:

1.    Familiarization EVA

2.    Communication ‘Lego Bricks’ experiment

3.    Sleep Study / crew medical questionnaires

4.    Musk Observatory Training and Observation

5.    Journalist interviews and filming

6.    Medical status surveys with Crew Doctor throughout the day.

7.    CPR Training

a.    Practice CPR techniques with crew medical doctor on CPR medical dummy

See schedule for tomorrow attached to this email.

Anomalies in work:

8.    Non Functional WiFi Modem

9.    Non Functional Power Strip (removed and replaced)

10. 1 Non Functional Radio

11. EVA Airlock Outer Hatch Failure (failed open) due to high winds.

12. Roof Hatch blown open continually (will close when winds die down)

Weather:  Very windy!  We have had two wind related MDRS failures.

Crew Physical Status:  All crew in good health.

EVA: No EVA today

Reports to be filed:

Greenhab

Operations

EVA Request

Journalist

Sol Summary

Mission Plan

Support Requested:

1.    We can’t find a template for Daily HSO Report.  Please provide what is desired in this report.

2.    Water Transfer from Delivery Tank to Static Tank

3.    More TB-Cide Quat Cleaner (we are out)

4.    Recommendation for wi-fi router failure.

a.    We have reset the router, checked our allocation, restarted laptop, and router is non-functional.  When hooking up laptop directly to modem internet works well.

Greenhab Report – December 3rd

GreenHab Report

Trisha Randazzo

December 3, 2017

Sol 2

Environmental control: (Choose which is appropriate and explain further if needed)

Ambient (no heating or/cooling)

Ambient with window/door open

· Windy so unable to keep door open

Heating

· Functioning nominally

Cooling

· Not availabloe

Both heating and cooling

· Nominal

Shade cloth on/off

· On

Average temperature: (N/A until new sensor is delivered)

· N/A

Changes to crops: Note all emergence of seedlings, death of seedlings, etc

· One tomato is doing poorly, but everyone else is growing quite well!!

Daily water usage for crops:

· Complete with no issues

Time(s) of watering for crops:

· 12:55

Morning research observations: (Currently not operational)

· N/A

Changes to research plants:

· N/A

Daily watering and amount of water used:

· 2 gallons of all soils

Aquaponics: (Currently not operational)

· N/A

Narrative: Any other information you want to share

· Planted crew gifts of paper-whites. Two containers in the Green-Hab, one container in the Hab.

Support/supplies needed:

· Awaiting: